Solve by completing the square:10x=3x^2-1
Answer by stanbon(75887)

Solve by completing the square:10x=3x^2-1
---------------------------------------------
3x^2 -10x -1 = 0
----
3(x^2- (10/3)x + ?) = 1/3 + 3?
----
3(x^2 - (10/3)x + (5/3)^2) = (1/3) + 3(5/3)^2
-----
3(x-(5/3)^2 = (1/3) + (25/3)
---
(x-(5/3))^2 = (26/9)
----
x-(5/3) = sqrt(26)/3 or x - (5/3) = -sqrt(26)
--------
x = (-5+sqrt(26))/3 or x = (5-sqrt(sqrt(26))/3
--------------------
Cheers,
Stan H.